# PALM: Property Attestation for Large Generative Models

### v1

@@ 4.2 Dealing with Large Datasets
To accurately attest to properties that depend on a dataset, PALM must measure the dataset (or its items) upon use, since $\sf{\mathcal{A}dv}$ could tamper with data in external storage at any time (recall threat model in Section 3.2). Figure 4 shows how PALM handles this for in-memory and memory-mapped datasets.
Case 1: In-memory Dataset. For integrity measurement, this is the ideal approach when sufficient TD memory is available. Since the TD memory is protected, data within cannot be tampered. As a result, the dataset can be measured once during loading, as long as it stays in memory for the entire operation. As shown in Figure 4 (left), our framework loads the entire dataset into memory, measures it, and has it available for the program thereafter. Since one hash operation is required, the overhead is minimal (discussed further in Section 6).
@@
Compared to straightforward dataset hashing, this approach incurs additional overhead due to iterative hash and modular operations. However, this overhead is negligible compared to computation-heavy ML tasks (shown in Section 6).
@@ 5.2 Metrics
  • Baseline: The baseline represents running the task on a TDX-enabled machine without any measurement or attestation from PALM. We omit the baseline for running our framework on non-TDX machines because the overhead of TDX compared to non-TDX systems has already been studied [confidential_llm_tdx_h100]. Our focus is solely on the overhead introduced by our framework in the environment where it is intended to be used.
@@ 6.1 R1: Efficient
The actual computation for these dataset-related tasks is not expensive compared to training and evaluation tasks. Thus, the observed 62% to 70% reflects the raw overhead from hashing, since the baseline workloads are minimal for calculating an attribute distribution and preprocessing. As for binding, the main overhead comes from iterating through the dataset to compute the MSH for all records. The input measurement time for iterating through the dataset is similar to the proof of preprocessing experiment, demonstrating consistent measurement time when the dataset remains the same.
@@
As for model optimization, similar to training, proof of fine-tuning (Table 4) also shows minimal overhead of $\leq$1.35% across all models, plus a substantial reduction in memory usage from 85-87 MB to only 4 MB. For proof of quantization, since no dataset is involved, the 4.7% overhead comes from measuring models.
@@ 8 Discussions
Other Data Integrity Techniques. PALM uses MSH to reduce datasets to a single value that can be computed incrementally as each item is read from storage, yielding the same result irrespective of the order in which the records are used. Nevertheless, the MSH is not without limitations. While the ordering of accesses does not impact the final hash, it is necessary to access each item exactly once, leading to unnecessary data accesses whenever a metric does not access the entire dataset. An alternative approach is to use a Merkle tree that allows arbitrary random access within a dataset, at the cost of $\mathcal{O}(\operatorname{log}n)$ additional reads and hash evaluations per lookup. An implementation might use a proof of binding to bind a dataset hash to both an MSH and a Merkle tree root hash, allowing it to select a hash type according to the access patterns of the metric being computed.
Selecting Chal in Property Attestation. Section 4.4 outlines a property attestation protocol enabled by PALM initiated by request from $\sf{\mathsf{Inr}}$ containing Chal. In conventional RA settings, Chal is a nonce or counter used to prevent replay attacks and ensure response freshness. Given ML property attestation in this work is non-interactive with $\sf{Vrf}$, selection of Chal should be derived from a timestamp rather than a traditional cryptographic nonce or counter.

### v2

@@ 4.2 Dealing with Large Datasets
To accurately attest to properties that depend on a dataset, PALM must measure the dataset (or its items) upon use, since $\sf{\mathcal{A}dv}$ could tamper with data in external storage at any time (recall Section 3.2). Figure 4 shows how PALM handles this for in-memory and memory-mapped datasets.
Case 1: In-memory Dataset. For integrity measurement, this is the ideal approach when sufficient TD memory is available. Since the TD memory is protected, data within cannot be tampered. As a result, the dataset can be measured once during loading, as long as it stays in memory for the entire operation. As shown in Figure 4 (left), our framework loads the entire dataset into memory, measures it, and uses it with the operation thereafter. Since one hash operation is required, the overhead is minimal (discussed further in Section 6).
@@
Compared to straightforward dataset hashing, this approach incurs additional overhead due to iterative hash and modular operations. However, it ensures no other countermeasures are required to account for random sampling, as any permutation of the full dataset will result in the same final MSH. Furthermore, this overhead is negligible compared to computation-heavy ML tasks (shown in Section 6).
@@ 5.2 Metrics
  • Baseline: The baseline represents running the task on a TDX-enabled machine without any measurement or attestation from PALM. We omit the baseline for running our framework on non-TDX machines because the overhead of TDX compared to non-TDX systems has already been studied [confidential_llm_tdx_h100].
@@ 6.1 R1: Efficient
The actual computation for these dataset-related tasks is not expensive compared to training and evaluation tasks. Thus, the observed 62.55% to 69.55% reflects the raw overhead from hashing, since the baseline workloads are minimal for calculating an attribute distribution and preprocessing. As for binding, the main overhead comes from iterating through the dataset to compute the MSH for all records. The input measurement time for iterating through the dataset is similar to the proof of preprocessing experiment, demonstrating consistent measurement time when the dataset remains the same.
@@
As for model optimization, proof of fine-tuning (Table 4) also shows minimal overhead of $\leq$1.35% across all models, plus a substantial reduction in memory usage from 85-87 MB to only 4 MB. For proof of quantization, since no dataset is involved, the 4.7% overhead comes from measuring models.
@@ 8 Discussions
Other Data Integrity Techniques. PALM uses MSH to reduce datasets to a single value that can be computed incrementally as each item is read from storage, yielding the same result irrespective of the order in which the records are used. Nevertheless, MSH is not without limitations. While the ordering of accesses does not impact the final hash, it requires that every record be accessed exactly once. As a result, MSH is only applicable to operations that use the entire dataset. Such cases are relatively common for LLMs, which typically require large amounts of data. If partial access within a dataset is required, an alternative approach is to use a Merkle tree, at the cost of $\mathcal{O}(\log n)$ additional reads and hash evaluations per lookup. An implementation might use a proof of binding to bind a dataset hash to both an MSH and a Merkle tree root hash, allowing it to select a hash type based on whether an operation requires access to the entire dataset.
Selecting Chal in Property Attestation. Section 4.4 outlines a property attestation protocol enabled by PALM initiated by request from $\sf{\mathsf{Inr}}$ containing Chal. In conventional remote attestation settings, Chal is a nonce or counter used to prevent replay attacks and ensure response freshness. Given ML property attestation in this work is non-interactive with $\sf{Vrf}$, selection of Chal should be derived from a timestamp rather than a traditional cryptographic nonce or counter.